Assume that when adults with smart phones are randomly selected 44% of them use them in meetings or classes if 12 adult smartphone users are randomly selected find the probability that fewer than 4 of them use their smart phones in meetings or classes


1
Expert's answer
2022-04-21T15:55:27-0400

This is a binomial distribution with n12, p=0.44.

"P(X<4)=P(X=0)+P(X=1)+P(X=2)+P(X=3)="

"=C_{12}^00.44^0(1-0.44)^{12-0}+C_{12}^10.44^1(1-0.44)^{12-1}+C_{12}^20.44^2(1-0.44)^{12-2}+C_{12}^30.44^3(1-0.44)^{12-3}=0.1502."